OpenStudy (anonymous):

Compute the value of the discriminant and give the number of real solutions of the quadratic equation.

OpenStudy (uri):

discriminant b-4ac

OpenStudy (solomonzelman):

the discriminant is, \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \LARGE{x=~} \huge{ \color{magenta}{b} ^2-4 \color{blue}{a} \color{red}{c}}~ }}}\) when the equation is \(\LARGE\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\).

OpenStudy (anonymous):

\[-4x^2+8x-1\]

OpenStudy (anonymous):

this is the equation

OpenStudy (uri):

a=-4 b=8 c=-1 plug this in =b-4ac

OpenStudy (anonymous):

so

OpenStudy (anonymous):

\[-8\pm \sqrt{64-16}\]

OpenStudy (anonymous):

I am not sure

OpenStudy (solomonzelman):

\(\bbox[8pt, lime,border:8px solid black]{\LARGE \LARGE{x=~} \huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{b} ^2-4 \color{blue}{a} \color{red}{c}}}{2 \color{blue}{a}} } ~ }\) when the equation is \(\LARGE\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\). (I think you are missing some part of the quadratic formula)

OpenStudy (anonymous):

oh all over -8

OpenStudy (solomonzelman):

yes

OpenStudy (uri):

Yeah. Now can you solve it further?

OpenStudy (anonymous):

yeah

OpenStudy (uri):

Alright. :)

OpenStudy (anonymous):

\[\frac{ -2\pm \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

@uri @SolomonZelman

OpenStudy (anonymous):

@uri @SolomonZelman

OpenStudy (anonymous):

@uri @SolomonZelman

OpenStudy (uri):

Must be right!

OpenStudy (anonymous):

so then what is the value of the descriminant and the number of real solutions

OpenStudy (anonymous):

@Nnesha @SolomonZelman @TheSmartOne

OpenStudy (accessdenied):

The value of the discriminant is simply the value under the square root in the quadratic formula. Here I will call it \( \Delta \) (because that is what I am used to labelling it as; others might use d or D instead). \( x = \dfrac{-b \pm \sqrt{\color{goldenrod}{b^2 - 4ac}}}{2a} = \dfrac{-b \pm \sqrt{\color{goldenrod}\Delta}}{2a} \) \( \begin{align} \Delta &= b^2 - 4ac \\ &= (8)^2 - 4(-4)(-1) \\ &= 64 - 16 \\ &= 48 \end{align} \) The discriminant tells you the number of real solutions based on its sign value. Think about how the quadratic formula is affected by the value there. If it is negative, the square root of a negative is not a real number! \( \Delta > 0 \), the quadratic has two unique real solutions \( \Delta = 0 \), the quadratic has one repeated solution (so, there are two but they are the same). \( \Delta < 0 \), the quadratic has no real solutions (two complex solutions)

OpenStudy (anonymous):

help me is the discriment \[\frac{ \pm 1 \sqrt{3} }{ -2 }\]

OpenStudy (accessdenied):

Here, the discriminant's value is 48 (remember, \(b^2 - 4ac \) is the discriminant), so the sign is positive. Thus, the solutions would be as indicated above.

OpenStudy (anonymous):

so the descriminant is 48 and there are 2 unique real solutions

OpenStudy (accessdenied):

Yup. b^2 - 4ac = (8)^2 - 4(-4)(-1) = 64 - 16 = 48. 48 > 0, so sqrt(48) is a real number and there are two real number solutions to the quadratic!

OpenStudy (anonymous):

ok thanks yay

OpenStudy (accessdenied):

No problem! :)

OpenStudy (uri):

Sorry,I was away but @AccessDenied helped. cx

OpenStudy (anonymous):

thx anyways ;)